OpenStudy (anonymous):

In a certain population, 15% of the people have Rh-negative blood. A blood bank serving this population receives 92 blood donors on a particular day. a)What is the probability that more than 80 of the donors are Rh-negative? someone please help. I get a z score of 0

OpenStudy (aaronq):

I think you can use the binomial distribution, find \(P(x>80)=1-P(x\leq80)\)

OpenStudy (anonymous):

ok. umm could you help with another question

OpenStudy (aaronq):

i may be able to, may not though. post it

OpenStudy (anonymous):

Suppose a random sample of n =25 observations is selected from a population that is normally distributed, with mean equal to 106 and standard deviation equal to 12. (a)Find the probability that the sample mean deviates from the population mean μ = 106 by no more than 4.

OpenStudy (aaronq):

okay so find \(P(106\pm4)\) \(z=\dfrac{110-106 }{12}=0.333\) \(P(z=0.333)=0.1293\) (from a z-table) because it's symmetric =0.1293*2=0.2586

OpenStudy (anonymous):

oh ok. thank you

OpenStudy (aaronq):

I'm not entirely sure this is what they're asking for, what i did is basically the probability of sampling within \(\pm4\) of the mean.  But i think it's right

OpenStudy (anonymous):

yeah.....it makes sense. thank you.

OpenStudy (aaronq):

okay cool, no problem !

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

how is the z score for 0.333 not 0.6293

OpenStudy (aaronq):

The z-table i have gives values like this